Locate and classify all critical points for f(x, y) = 4y^2x + 3xy


1
Expert's answer
2022-02-21T16:12:26-0500

Find the critical point(s)


"\dfrac{\partial f}{\partial x}=4y^2+3y"

"\dfrac{\partial f}{\partial y}=8xy+3x"

Then


"\begin{cases}
 \dfrac{\partial f}{\partial x}=0 \\ \\
 \dfrac{\partial f}{\partial y}=0
\end{cases}"

"\begin{cases}
 4y^2+3y=0 \\ 
 8xy+3x=0
\end{cases}"

Critical points "(0, 0), (0, -3\/4)"

Point "(0,0)"


"\dfrac{\partial^2 f}{\partial x^2}(0,0)=0"

"\dfrac{\partial^2 f}{\partial y^2}(0,0)=8(0)=0"

"\dfrac{\partial^2 f}{\partial x\partial y}(0,0)=\dfrac{\partial^2 f}{\partial y\partial x}(0,0)=8(0)+3=3"


"D(0,0)=\begin{vmatrix}
 0 & 3 \\
 3 & 0
\end{vmatrix}=-9<0"

Then the "f(0,0)" is not a local maximum or minimum.

The critical point "(0,0)" is a saddle point.


Point "(0,-3\/4)"


"\dfrac{\partial^2 f}{\partial x^2}(0,-3/4)=0"

"\dfrac{\partial^2 f}{\partial y^2}(0,-3/4)=8(0)=0"

"\dfrac{\partial^2 f}{\partial x\partial y}(0,-3/4)=\dfrac{\partial^2 f}{\partial y\partial x}(0,-3/4)"

"=8(-3\/4)+3=-3"


"D(0,0)=\begin{vmatrix}
 0 & -3 \\
 -3 & 0
\end{vmatrix}=-9<0"

Then the "f(0,-3\/4)" is not a local maximum or minimum.

The critical point "(0,-3\/4)" is a saddle point.